### v1

# Discovering Concept Directions from Diffusion-based Counterfactuals via Latent Clustering

Abstract
Concept-based explanations have emerged as an effective approach within Explainable Artificial Intelligence, enabling interpretable insights by aligning model decisions with human-understandable concepts. However, existing methods rely on computationally intensive procedures and struggle to efficiently capture complex, semantic concepts. Recently, the Concept Discovery through Latent Diffusion-based Counterfactual Trajectories (CDCT) framework, introduced by Varshney et al. (2025), attempts to identify concepts via dimension-wise traversal of the latent space of a Variational Autoencoder trained on counterfactual trajectories. Extending the CDCT framework, this work introduces Concept Directions via Latent Clustering (CDLC), which extracts global, class-specific concept directions by clustering latent difference vectors derived from factual and diffusion-generated counterfactual image pairs. CDLC substantially reduces computational complexity by eliminating the exhaustive latent dimension traversal required in CDCT and enables the extraction of multidimensional semantic concepts encoded across the latent dimensions. This approach is validated on a real-world skin lesion dataset, demonstrating that the extracted concept directions align with clinically recognized dermoscopic features and, in some cases, reveal dataset-specific biases or unknown biomarkers. These results highlight that CDLC is interpretable, scalable, and applicable across high-stakes domains and diverse data modalities.
keywords:
Explainability , Counterfactual Explanation , Concept-Based Explanation , Latent Diffusion Model , Dermoscopy , Concept Directions , Clustering
journal: Pattern Recognition Letters
[…]
1 Introduction
In high-stakes applications, such as medical diagnosis, financial risk assessment, and autonomous driving, understanding the rationale behind a neural network’s decision is often as important as the decision itself. Explainable Artificial Intelligence (XAI) [1, 2] has therefore emerged as a critical research area, aiming to bridge the gap between high-performing black-box models and human interpretability. Among the various XAI paradigms, concept-based explanations [3, 4] have gained particular attention due to their ability to express model behavior in terms of high-level, semantically meaningful concepts, rather than low‑level feature weights or pixel-based saliency maps [5, 6]. By aligning explanations with concepts recognized by domain experts, these methods facilitate trust [7, 8], debugging [9], and regulatory compliance [10, 11].
Although concept-based explainability has been widely explored using convolutional [3, 4] and GAN-based architectures [12, 13], its application within diffusion-based generative models remains relatively underexplored. Recent works, such as Concept-Guided Latent Diffusion Counterfactual Explanations (CoLa-DCE) [14], have demonstrated the ability to produce spatially constrained, concept-conditioned counterfactuals for an arbitrary classifier. However, this approach only utilizes concepts to guide counterfactual explanations and does not extract global concept representations that are generalizable across examples.
To discover global concepts that generalize across multiple samples, Varshney et al. [15] proposed the Concept Discovery through Latent Diffusion-based Counterfactual Trajectories (CDCT) framework. CDCT generates a classifier-guided counterfactual image trajectory dataset using a latent diffusion model [16] and subsequently trains a Variational Autoencoder (VAE) [17] on this trajectory dataset to disentangle classifier-relevant concepts. While CDCT represents a significant advancement in leveraging diffusion-based counterfactuals for concept discovery, it relies on a dimension-wise traversal strategy, wherein each latent variable is modified independently to detect relevant concepts. This exhaustive search procedure is computationally expensive, particularly in high-dimensional latent spaces, and inherently overlooks semantic concepts that arise from interactions among multiple latent dimensions. Consequently, CDCT often identifies only simple concepts, restricting its capacity to uncover high-level semantic directions.
To address these limitations, this paper proposes a novel framework Concept Directions via Latent Clustering (CDLC), that extracts multi-dimensional concepts by clustering latent difference vectors computed from factual–counterfactual image pairs. These difference vectors, derived from VAE encodings of factual images and their corresponding diffusion-based counterfactuals, capture the classifier-induced transformation in latent space. In contrast to CDCT, which modifies individual latent dimensions, CDLC leverages directional clustering to reveal coordinated latent changes that correspond to semantically meaningful concepts. This approach not only reduces computational complexity by eliminating exhaustive per-dimension search but also enables the discovery of classifier-relevant concept directions that emerge from interactions among multiple latent dimensions, effectively overcoming a limitation of the original CDCT formulation.
This paper makes the following key contributions:
  • 1.
    A novel framework, Concept Directions via Latent Clustering (CDLC), is introduced as an extension of CDCT [15], significantly reducing computational complexity compared to dimension-wise search.
  • 2.
    CDLC extracts global, multidimensional semantic concept directions by clustering latent-difference vectors obtained from VAE encodings of factual and diffusion-generated counterfactual images.
  • 3.
    The effectiveness of CDLC is validated on a skin lesion classification task, where the discovered concept directions not only reliably flipped classifier predictions but also transferred robustly to unseen samples.
2 Related Work
Explainable Artificial Intelligence (XAI) aims to make machine learning models more transparent by providing human-understandable insights into their decision-making processes [1, 2]. Among various XAI approaches, concept-based explanations have gained increasing attention due to their ability to link internal model representations to semantically meaningful, human-interpretable concepts [3, 4, 18]. In supervised settings, concept-based methods rely on expert-provided annotations, which are costly and time-consuming to collect [3, 19]. For instance, Testing with Concept Activation Vectors (TCAV) [3] requires example sets for each concept and learns linear concept activation vectors to quantify a model’s sensitivity to those concepts. While effective, this reliance on curated concept examples limits the scalability and generalization of such methods. To overcome these limitations, unsupervised concept-based methods have been introduced [4, 20]. A prominent example is Automatic Concept Explanations (ACE) [4], which discovers concepts by clustering segmented image patches, without requiring manual annotation. However, ACE relies on spatial assumptions by considering localized image regions as potential concept candidates, limiting its ability to capture non-local or abstract concepts. A comprehensive survey [18] has reviewed the landscape of concept-based explanations, outlining a range of methodologies and their respective applications.
Another prominent paradigm in XAI is counterfactual explanations, which provide “what-if” scenarios to reveal how minimal semantic changes to an input can alter a model’s prediction [21]. With the emergence of generative models, counterfactual synthesis has advanced significantly, enabling the generation of more realistic and semantically meaningful examples.
In particular, diffusion-based approaches have been proposed to generate counterfactuals by incorporating classifier gradients into the denoising process [22, 23] or by applying adaptive parameterization and cone regularization of gradients [24]. While these methods produce visually realistic counterfactuals, their explanations are typically local to individual samples. In contrast, the Global Counterfactual Directions (GCD) [25] method learns latent directions that consistently invert classifications across samples via a proxy model, enabling more generalizable interpretability.
Recent work has attempted to combine concepts with counterfactual generation. CoLa‑DCE [14] uses classifier guidance to produce spatially constrained, concept‑conditioned counterfactuals. However, these edits remain sample-specific and do not yield reusable global concept vectors. In contrast, Varshney et al. [15] leveraged classifier-guided counterfactual trajectories to identify global disentangled semantic concepts. Despite its strengths, CDCT’s dimension-wise traversal is computationally expensive and limited in detecting complex, multidimensional concepts.
Another line of research leverages clustering in latent spaces to model semantic variation. For example, DifCluE [26] clusters latent embeddings produced by a diffusion autoencoder to generate diverse counterfactual explanations, effectively modeling intra-class variation, but does not explicitly identify global concept directions.
However, most existing approaches reveal several open challenges: reliance on concept annotations or spatial heuristics, a focus on localized or sample-specific edits, the need for exhaustive latent traversal, and limited capacity to capture global, multidimensional semantic concepts. To address these limitations, a novel framework, Concept Directions via Latent Clustering, is proposed as an extension of CDCT [15]. It discovers global, class-specific concept directions by clustering directional latent differences between factual and counterfactual images. This approach enables the extraction of complex, multidimensional concept directions in unsupervised manner and offers a computationally efficient framework for concept-based explanation, enhancing both interpretability and scalability.
3 Methodology
Figure 1: Overview of the CDLC framework. Counterfactuals are generated using a Latent Diffusion Model (LDM) with classifier guidance, following the procedure used in CDCT [15]. Factual–counterfactual image pairs are encoded using a pretrained Variational Autoencoder (VAE), and the difference between their latent representations is normalized to form unit vectors. These vectors are clustered to identify class-specific concept directions $C_{j}$. During inference, each direction, scaled by a factor $\alpha$, is applied to the test sample’s latent representation to observe its effect on classifier output.
Concept Directions via Latent Clustering (CDLC) is a novel framework designed to extract global, multidimensional concept directions from factual and counterfactual image pairs. It builds upon the counterfactual generation stage of CDCT [15], but instead of relying on computationally expensive dimension-wise latent traversals, it identifies concept directions by clustering unit-normalized difference vectors in a semantic latent space. This section describes the key components of CDLC in detail.
3.1 Counterfactual Generation via Latent Diffusion Models
Given an input image $x_{f}$ and a trained classifier $f(\cdot)$, a counterfactual image $x_{cf}$ is synthesized to belong to a target class $y_{cf}\neq f(x_{f})$. The counterfactual generation process follows the first stage of the CDCT framework [15], termed Generation of Counterfactual Trajectories, where classifier guidance is integrated into the denoising steps of a latent diffusion model [16] to produce semantically meaningful modifications aligned with the desired class transition. Details of this process can be found in the CDCT framework [15]. Rather than capturing intermediate reconstructions across the trajectory, only the final counterfactual image is retained. This design choice aims to capture the complete semantic shift required to alter the classifier’s prediction while also reducing storage and computational overhead. The resulting counterfactual is visually realistic and deviates minimally from the factual image, thereby isolating class-specific semantic changes.
3.2 Discovery of Concept Directions
Each factual image $x_{f}$ and its corresponding counterfactual $x_{cf}$ for a specific target class are encoded using a pretrained Variational Autoencoder, resulting in latent representations $z_{f}=\mathrm{VAE}(x_{f})$ and $z_{cf}=\mathrm{VAE}(x_{cf})$. These latent embeddings capture high-level semantic characteristics of the original and counterfactual images in a compact form.
A difference vector is computed between the latent representations of each factual and counterfactual image pair as:
$\Delta z=z_{cf}-z_{f}$
Each difference vector is reshaped into a flat vector $\Delta z\in\mathbb{R}^{d}$. To focus solely on the directional semantics and eliminate the influence of vector magnitude, each difference vector is normalized to unit length:
$\tilde{z}=\frac{\Delta z}{\|\Delta z\|_{2}}$
This transformation projects the latent differences onto the unit hypersphere in $\mathbb{R}^{d}$, making them suitable for angular similarity-based clustering.
A collection of such unit-norm latent difference vectors $\{\tilde{z}_{i}\}_{i=1}^{N}$, derived from multiple factual–counterfactual pairs for a given target class, is clustered into $K$ clusters to identify shared semantic transformations. Cluster centers $C=\{c_{1},c_{2},\ldots,c_{K}\}$ are computed by averaging the unit vectors within each cluster and re-normalizing to unit length:
$c_{k}\in\mathbb{R}^{d},\quad\|c_{k}\|_{2}=1$
Each cluster center $c_{k}$ represents a global concept direction, a consistent and interpretable shift in the latent space associated with the classifier’s decision boundary. These concept directions can then be applied to the latent representations of new test samples to induce semantically meaningful changes in the reconstructed output.
3.3 Concept Interpretation
To interpret each discovered concept direction $c_{k}$, it is applied to the latent encoding of an unseen test sample:
$z^{\prime}_{\text{test}}=z_{\text{test}}+\alpha\cdot c_{k},$
where $\alpha\in\mathbb{R}$ controls the strength of the semantic manipulation. The modified latent vector $z^{\prime}_{\text{test}}$ is then passed through the VAE decoder to generate a concept-modified image.
In contrast to CDCT, which iteratively modifies each latent dimension to search for influential concepts, CDLC learns global, multi-dimensional concept directions directly from the difference of factual–counterfactual encodings. This leads to a significant increase in computational efficiency and interpretable semantic transformations.
4 Experiments & Results
This section presents the experimental setup and an analysis of the results produced by Concept Directions via Latent Clustering framework. Both quantitative and qualitative evaluations are performed to assess the semantic coherence of the extracted concept directions and their effectiveness in influencing classifier predictions.
4.1 Dataset and Classification Model
The same dataset and classification architecture as used in the CDCT framework [15] are adopted for evaluation. Experiments are conducted on a consolidated dermoscopic image dataset derived from the International Skin Imaging Collaboration (ISIC) challenges (2016–2020) [^1] Dataset available at: https://challenge.isic-archive.com/challenges/.. Following the duplicate removal strategy proposed by Cassidy et al. [27], a curated dataset of 29,468 unique images is obtained and stratified into training, validation, and test subsets. The dataset includes eight diagnostic categories: Melanocytic Nevus (NV), Melanoma (MEL), Basal Cell Carcinoma (BCC), Actinic Keratosis (AK), Benign Keratosis (BKL), Dermatofibroma (DF), Vascular Lesions (VASC), and Squamous Cell Carcinoma (SCC).
A ResNet-50 [28] model is trained on the ISIC training partition and used as the target classifier throughout all experiments. This network is used for both generating classifier-guided counterfactuals and evaluating the effectiveness of discovered concept directions.
4.2 Generation of Counterfactuals
The same classifier-guided counterfactual generation approach introduced in the CDCT framework [15] is adopted, with a simplified setting: rather than capturing intermediate reconstructions across the trajectory, only the final counterfactual image is retained. Counterfactuals are generated using a Latent Diffusion Model based on the Stable Diffusion (SD) 2.1 architecture [^2] Model available at: https://huggingface.co/stabilityai/stable-diffusion-2-1., which is fine-tuned on the consolidated ISIC training dataset to align the generative manifold with the dermoscopic image domain. For each training sample, counterfactuals are synthesized for all target classes other than the class predicted by the classifier.
4.3 Extraction of Concept Directions
Semantic concept directions induced by classifier-guided counterfactuals are explored in two distinct latent spaces: (1) the pretrained encoder of the Stable Diffusion 2.1 model, referred to as the LDM encoder, and (2) a Variational Autoencoder trained on counterfactual trajectories, referred to as the CDCT encoder (see Step 2 of the CDCT framework [15] for further details). Experimental setup, results, and analysis based on the CDCT encoder are provided in the supplementary material.
To enable class-specific concept discovery, latent difference vectors are computed separately for each target class. For this purpose, training samples not predicted as the target class are selected, and counterfactual images are generated that shift the classifier’s prediction to the target class. The corresponding factual and counterfactual images are encoded using the LDM encoder, resulting in latent embeddings of shape $4\times 32\times 32$. The element-wise difference between these embeddings captures the transformation required to alter the prediction. Each difference tensor is flattened into a $4096$-dimensional vector and normalized to unit length, forming a directional vector in latent space.
The unit vectors are collected for all training samples that are not predicted as a given target class. Spherical K-Means clustering [29] is then applied to group these vectors that reflect similar semantic transformations. The number of clusters is selected based on the silhouette coefficient to ensure cohesion and separation. Each resulting cluster represents a set of consistent latent transformations, and the average unit direction within a cluster is interpreted as a representative concept direction for the target class.
4.4 Interpretation and Evaluation of Concept Directions
(a) Success Rate = 70.54 % - Purplish Core Pigmentation.
(b) Success Rate = 83.61 % - Reddish Core with Blue-Gray Dots.
(c) Success Rate = 81.09 % - Blotchy Pigmentation with Irregular Texture.
(d) Success Rate = 85.44 % - Central Pinkish Veil with Asymmetric Intensified Pigmentation.
(e) Success Rate = 84.59 % - Central Purplish Veil.
(f) Success Rate = 70.47 % - White Structures with Irregular Vessels.
Figure 2: Discovered concepts by CDLC on the ISIC dataset using the LDM encoder. Each row shows two examples: original, reconstructed, and manipulated reconstruction (left to right). The predicted probability for the target class associated with each concept direction is shown above each image.
To assess the interpretability and effectiveness of the extracted concept directions, each identified concept direction is applied to unseen test images in the latent space. Figure 2 illustrates concept directions identified by CDLC using the LDM encoder. The success rate reported in each sub-caption indicates the proportion of test samples for which traversal along the corresponding direction increases the predicted probability of the target class, relative to the reconstructed image.
For Melanoma, the extracted direction (Figure 2a) increases central pigmentation and introduces a purplish hue in the core lesion. In Basal Cell Carcinoma, detected concept direction (Figure 2b) produces a reddish-purple hue with blue-gray dots. For Benign Keratosis, a concept direction (Figure 2c) induces purplish blotches, irregular pigmentation, and blurred borders. In the case of Dermatofibroma, the direction (Figure 2d) introduces a violet-toned center with subtle border irregularities and increased asymmetric pigmentation. For Vascular Lesions, the direction (Figure 2e) enhances central coloration with purplish tones and peripheral pink diffusion. Finally, Squamous Cell Carcinoma directions (Figure 2f) reveal white central structures accompanied by irregular linear streaks suggestive of vascular features.
5 Discussion
The results demonstrate that CDLC effectively extracts global concept directions for skin lesion classification while significantly improving computational efficiency over CDCT. By avoiding iterative traversal of individual latent dimensions, CDLC reduces the overhead typically associated with concept discovery, especially in high-dimensional latent spaces. Additionally, it captures rich, multidimensional semantic directions beyond the axis-aligned perturbations of CDCT, which tend to detect only simple features. Beyond efficiency, CDLC reliably discovers class-specific concept directions by constructing latent difference vectors separately for each target class. This avoids feature entanglement observed in CDCT, where a single VAE is trained on trajectories across all classes. As a result, the extracted directions are more semantically coherent and clinically interpretable.
To further analyze the visual characteristics of the extracted directions, the behavior of the two latent encoders used in the framework is compared. While both encoders support meaningful concept discovery, the CDCT encoder often fails to reconstruct subtle yet clinically relevant features, limiting the fidelity of concept manipulations. In contrast, the LDM encoder preserves fine-grained details and structural consistency, yielding more realistic and diagnostically coherent manipulations. Nonetheless, directions in the LDM space frequently introduce pinkish or violet hues, possibly reflecting biases from the generative prior, which may influence the visual appearance of concepts.
The concept directions discovered by CDLC demonstrate alignment with established dermoscopic features. For instance, directions targeting Melanoma often exhibit asymmetric pigmentation and color variegation, features of malignancy reported in clinical literature [30]. In Basal Cell Carcinoma, the extracted directions introduce translucent reddish or violet hues with dark pigment patches, consistent with features such as ulceration and blue-gray dots [31]. For Dermatofibroma, directions emphasize a central scar-like area of pink hues reported in literature [32], while Vascular Lesions display purplish centers with surrounding pink diffusion, resembling the vascular blush of angiomas [33]. Similarly, directions for Squamous Cell Carcinoma reveal white scaly textures and irregular vascular patterns, characteristic of keratinizing and sun-damaged lesions [34].
In addition to replicating known diagnostic cues, CDLC also uncovers subtler variations such as peripheral pigmentation or central hue changes, that may indicate dataset-specific biases or underexplored clinical markers, requires clinical feedback. In contrast, reliable concept directions could not be identified for the Nevus class. This is likely due to class imbalance, as the dominance of Nevus in the training data results in fewer samples being mapped as counterfactual to it. Consequently, the limited number of latent difference vectors hinders the discovery of consistent semantic directions for this class.
6 Limitations & Future work
While CDLC demonstrates promising results in extracting semantically meaningful and class-specific concept directions, several limitations remain. The reliability of extracted directions depends on the quality of generated counterfactuals, artifacts or semantic drift in this step may introduce noise into the latent differences. Additionally, spherical clustering assumes that concept directions are well-separated based on angular similarity on a unit hypersphere, which may not adequately capture complex latent geometries or overlapping semantic structures.
Future work could explore alternative encoders, such as Diffusion Autoencoders [35], to enhance reconstruction fidelity and semantic expressiveness. Extending CDLC to other modalities, including multimodal or temporal data, may further broaden its applicability. Finally, incorporating human-in-the-loop evaluation or clinical feedback would support validation and real-world integration of discovered concepts.
7 Conclusion
This work introduced CDLC, a framework for discovering global, class-specific concept directions by clustering latent differences between factual and counterfactual image pairs. Experiments on a skin lesion classification task demonstrated that the extracted directions influence model predictions and align with known dermoscopic features, supporting their interpretability. However, some concept directions, especially obtained from the LDM encoder, consistently exhibit pinkish or purplish hues, irrespective of the target class. This visual redundancy may indicate an inherent bias in the generative prior, suggesting a need for further analysis to disentangling meaningful concepts from model or dataset-specific artifacts. The ability to uncover diverse, clinically relevant concepts positions CDLC as a valuable tool for advancing concept-based explainability in high-stakes domains.
8 Acknowledgments
The project was funded by the Federal Ministry for Education and Research (BMBF) with grant number 03ZU1202JA.
9 CRediT authorship contribution statement
Payal Varshney: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Software, Validation, Visualization, Writing – original draft. Adriano Lucieri: Conceptualization, Data curation, Supervision, Validation, Writing – review and editing. Christoph Balada: Writing – review and editing. Andreas Dengel: Funding acquisition, Project administration, Resources, Supervision, Writing – review and editing. Sheraz Ahmed: Supervision, Writing – review and editing.
10 Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used chatgpt in order to improve language and readability. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.
11 Data availability
This study uses publicly available dermoscopic images from the ISIC challenges (2016–2020). The exact dataset splits used for training, validation, and testing will be made available upon request.
Supplementary Material
Extraction of Concept Directions Using the CDCT Encoder
This supplementary section presents the experimental setup used to extract concept directions using the encoder from the CDCT framework (hereafter referred to as the CDCT encoder). The CDCT encoder is a Variational Autoencoder (VAE) trained on a counterfactual trajectory dataset generated via a latent diffusion model with classifier guidance, as described in Step 2 (Semantic Space Disentanglement) of the CDCT framework [15]. The training objective encourages the latent space to encode distinct, classifier-relevant transformations, facilitating more interpretable and disentangled concept representations. Further architectural and training details are provided in the original work [15].
In this supplementary analysis, the same pipeline described in Section 3.2 of the main manuscript is applied to compute latent difference vectors and extract concept directions. Specifically, for each training sample not predicted as the target class, a classifier-guided counterfactual is generated, and both the factual and counterfactual images are encoded using the CDCT encoder. Unlike the LDM encoder, which yields latent embeddings of shape $4\times 32\times 32$, the CDCT encoder produces compact latent representations of shape $512\times 1\times 1$. Each latent difference is flattened into a 512-dimensional vector and normalized to unit length to produce directional vectors. Spherical K-Means clustering [29] is then applied to the set of unit-norm latent differences computed per class. The number of clusters is selected based on the silhouette coefficient to capture cohesive and distinct semantic directions. The average unit direction within each resulting cluster is interpreted as a representative concept direction for the target class.
The results highlight how concept representations vary in the CDCT encoder’s latent space, particularly in terms of semantic disentanglement and reconstruction fidelity. Visualizations and class-specific examples are provided in the following subsections to support a comparative evaluation between the CDCT and LDM encoders.
Interpretation and Evaluation of Concept Directions Using the CDCT Encoder
To evaluate the semantic relevance and discriminative influence of the discovered concept directions, each direction identified by the CDCT encoder is applied to the latent representation of previously unseen test samples. Figure 3 presents representative concepts produced using the CDCT encoder. For each direction, the accompanying success rate (reported in each sub-caption) quantifies the proportion of test cases in which latent traversal increases the predicted probability of the corresponding target class, relative to the classifier’s output on the reconstructed factual image.
The concept directions extracted using the CDCT encoder exhibit class-specific semantic transformations aligned with known dermoscopic features. For Melanoma, the direction (Figure 3a) darkens the lesion periphery, reduces central contrast, and softens the borders. The Basal Cell Carcinoma direction (Figure 3b) reduces pigmentation and softens edge contrast, yielding a more blended lesion appearance. In the case of Actinic Keratosis (Figure 3c), the identified direction introduces a central lightened area with a scaly texture and peripheral reddish hue. The Benign Keratosis direction (Figure 3d) induces reddish central pigmentation with uneven borders and surrounding pigment variation. For Dermatofibroma, the concept direction (Figure 3e) enhances a pale central region surrounded by a subtle pink hue, features reported in literature [32]. The direction for Vascular Lesion (Figure 3f) reveals central redness accompanied and subtle irregularities in pigment distribution, resembling vascular blushes typically observed in angiomas and angiokeratomas [33] Finally, the Squamous Cell Carcinoma direction (Figure 3g) introduces a central white, keratin-like structure surrounded by inflamed reddish tissue, consistent with dermoscopic features described by Rosendahl et al. [34].
While several concept directions extracted using the CDCT encoder correspond to clinically relevant dermoscopic patterns, such as keratin-like centers in Squamous Cell Carcinoma or pale regions in Dermatofibroma, these findings are less detailed visually than those produced by the LDM encoder. The CDCT encoder, despite being explicitly trained to disentangle classifier-relevant transformations, demonstrates lower success rates across most classes in comparison to results produced by the LDM encoder. This can be attributed to its limited ability to reconstruct fine-grained dermoscopic features, which are critical for clinical interpretation. Furthermore, its lower-dimensional latent space restricts its capacity to model complex transformations, further limiting its effectiveness in capturing meaningful semantic changes.
Although many of the extracted directions align with features documented in dermatological literature, their diagnostic relevance remains to be clinically validated. Expert review by dermatologists is essential to determine whether these directions reflect genuine clinical cues, dataset-specific artifacts, or potentially novel biomarkers. Such validation would enhance the interpretability and utility of concept-based explanations in high-stakes medical applications.
(a) Success Rate = 65.62 % - Peripheral Darkening and Border Irregularity.
(b) Success Rate = 77.53 % - De-pigmentation and Soft-Border.
(c) Success Rate = 65.21 % - Scaly, Lightened Center with Red Halo.
(d) Success Rate = 70.03 % - Reddish Center with Surrounding Pigmentation.
(e) Success Rate = 76.17 % - Central Scar-Like Area with Pink Hue.
(f) Success Rate = 64.50 % - Central Redness.
(g) Success Rate = 85.64 % - White Center with Red Surround.
Figure 3: Discovered concepts by CDLC on the ISIC dataset using the CDCT encoder. Each row shows two examples: original, reconstructed, and manipulated images (left to right). The predicted probability for the target class associated with each concept direction is shown above each image.